OpenStudy (anonymous):

Helium is given in a temperature of 100 C. the average kinetic energy of helium is given by E=3/2*n*R*T. calculate the de broglie wavelength of the average helium atom in this temperature. if you can please write the whole calculation and not only the solution, tnx :)

OpenStudy (anonymous):

hey aaron, first of all (I'm just guessing by the name) if you're an Israeli we can switch to hebrew. secondly, I have all of the equations, I also have the final result. from some reason I'm not getting the same numbers that's why I asked for the way to solve it. if you can write it down that could be great .

OpenStudy (anonymous):

*for some reason

OpenStudy (aaronq):

not israeli, sorry. i mean it's just a question of plugging stuff in and solving, did you try it with both the relativistic and the non-?

OpenStudy (anonymous):

I calculated the average kinetic energy of one helium atom by using this formula: \[\epsilon= \frac{ 3 }{ 2 }*k_{b}*T\] then using the energy I found the average speed of the helium atom using this formula: \[E=\frac{ 1 }{ 2 }*m*V^{2}\] and then to find the de broglie wavelength of the average helium atom I used this formula: \[\lambda_{Db}=\frac{ h }{ p }=\frac{ h }{ m*v }\] for m I took the Mw of helium from the periodic table and divided it by Na (avogadro constant).

OpenStudy (anonymous):

\[\epsilon=7.7211\times 10^{-23}\] \[m=6.6456\times 10^{-24}\] v=48.2044 \[\lambda _{Db}=2.0665\times10^{-12}\] the right answer is \[\lambda _{Db}=65.5\times ^{-12}\]

OpenStudy (anonymous):

I found out that I forgot change unites for the mass and that I used gr instead of Kg so the mass should be 6.6456*10^-21

OpenStudy (anonymous):

but now the answer i get is 655e-12 instead of 65.5e-12

OpenStudy (aaronq):

theres definitely an error in the unit conversions your answer is way too close to not be that.

OpenStudy (anonymous):

k tnk you I'll keep on looking, it's good to know I'm on the right path :)

OpenStudy (aaronq):

no prob, good luck!